Verify Maclaurin's theorem for f(x)=(1-x)^(5/2) with Lagrange's form of remainder upto 3 terms where x=1
1
Expert's answer
2021-10-08T11:14:09-0400

"f(x)=f(0)+f'(0)x+{\frac {f''(0)x^{2}} 2} + {\frac {f'''(c)x^{3}} 6}"

where "c\in (0;x)""x=1"

f(1) = 0

f(0) = 1

"f'(x) = -{\frac 5 2} (1-x)^{{\frac 3 2}}"

f'(0) = -2.5

"f''(x) = {\frac {15} 4} (1-x)^{{\frac 1 2}}"

f''(0) = 3.75

f'''(c) = "-{\frac {15} {8\sqrt{1-c}}}"

So, we have to verify if "\exist c\in (0;1): 0=1-2.5+1.875-{\frac {15} {48\sqrt{1-c}}}"

"{\frac {15} {48\sqrt{1-c}}}=0.375"

"\sqrt{1-c} = {\frac 5 6}"

"c = {\frac {11} {36}}"

Theorem has been verified and proved.